# Stochastic Thermodynamics of Non-reciprocally Interacting Particles and Fields

### v2

@@ 4.4 Thermodynamic-Kinetic Uncertainty Relation
The thermodynamic uncertainty relation (TUR) delineates the trade-off between the precision of an observable current and the minimum EP required to sustain it [159, 234, 235, 236]. Here, the precision is the ratio of the square of the mean observable transition current and the variance of the transition current. TUR gives a tighter bound on the EP than the second law of thermodynamics. In other words, the precision of an observable current comes at a minimum EP required to sustain it. The bounds on the thermodynamics dissipation can be improved by choosing other appropriate physical observables and constraints [237, 238, 239, 240, 241, 242, 243, 244, 245, 246, 196, 238, 247, 248] and it’s unification to the thermodynamic-kinetic uncertainty relation (TKUR) [249]. The TUKR’s connection to the Speed Limits (SL) [250, 251, 252, 253, 254, 255, 256, 249, 166] and the FR symmetry [257, 258, 259, 260, 261] has been established. Here, we demonstrate the TKUR for non-reciprocal systems, a more fundamental analysis of the TKUR and non-quadratic speed limits is outlined in Ref.[88].
@@ A Coarse-grained observable transition currents
We consider a set of macroscopic observable transition currents denoted by $O^{J}=\{\Delta^{o}\}$ and the corresponding mean inferred EPR $\langle\dot{\Sigma}_{O^{J}}\rangle$. The mapping $\{\{\Delta_{\gamma\gamma^{\prime}}\},\{\Delta_{i}^{\vec{\mathcal{D}}}\}\}=\{\Delta\}\to\{\Delta^{o}\}$ is considered to be many to one, so that each microscopic transition contributes to one macroscopic observable transition current. This mapping is represented using an observable matrix $\mathbb{O}^{J}$ with the row and column index for the observable and microscopic current respectively. Mathematically, the mapping $\{\{\Delta_{\gamma\gamma^{\prime}}\},\{\Delta_{i}^{\vec{\mathcal{D}}}\}\}\to\{\Delta^{o}\}$ implies, $\mathbb{O}_{o^{\prime}i^{\prime}}^{J}=1\implies\mathbb{O}_{o^{\prime\prime}i^{\prime}}^{J}=0,\forall o^{\prime\prime}\in\{\Delta^{o}\}-o^{\prime}$. The observable current and the traffic vector satisfy $\vec{J}_{\Delta^{o}}=\mathbb{O}^{J}\vec{J}_{\Delta}$ and $\vec{T}_{\Delta^{o}}=\mathbb{O}^{J}\vec{T}_{\Delta}$. Notably, using eqs. 29 and 31 one can equivalently obtain the same expression for the mean and variance of the observable current. The mean inferred EPR $\langle\dot{\Sigma}_{O^{J}}\rangle$ reads [88]:
$\begin{split}\langle\dot{\Sigma}_{O^{J}}\rangle=\sum_{\{\Delta^{o}\}}2J_{\Delta^{o}}\tanh^{-1}{\left(\frac{J_{\Delta^{o}}}{T_{\Delta^{o}}}\right)},\end{split}$ (58)
$\langle\dot{\Sigma}\rangle\geq\langle\dot{\Sigma}_{O^{J}}\rangle$ holds true using the generalized log-normal inequality and the definition $\vec{J}_{\Delta^{o}}=\mathbb{O}^{J}\vec{J}_{\Delta}$ and $\vec{T}_{\Delta^{o}}=\mathbb{O}^{J}\vec{T}_{\Delta}$ [262]. The non-quadratic formulation gives a tighter bound on the EPR than the traditional-quadratic TUR [256] and is closely related to [263, 254, 249, 166]. The mean observable inferred EP $\langle{\Sigma}_{O^{J}}\rangle$ satisfies the inequality,
@@
[…]
@@ B Coarse-grained vorticities and state-correlations
Here, we aim to examine the EPR inference using state-space observables. We choose a specific sub-case of the set of vorticity currents between the observable macrostate $\rho_{o}=\mathbb{O}^{\rho}\vec{\rho}$, where $\mathbb{O}^{\rho}$ quantifies the many-to-one mapping from $\{\rho_{i}\}\to\{\rho_{o}\}$. $O^{\omega}=\{\omega_{oo^{\prime}}\}$ is quantified by the observable vorticity operator $\mathbb{O}^{\omega}$ that connects microscopic vorticities $\{\omega_{ij}\}$ to observable vorticities through $\vec{\omega}_{O}=\mathbb{O}^{\omega}\vec{\omega}$. The anti-symmetric (vorticity) and symmetric (traffic) are denoted by $\langle\omega_{oo^{\prime}}\rangle=\langle\rho_{o}\partial_{t}\rho_{o^{\prime}}-\rho_{o^{\prime}}\partial_{t}\rho_{o}\rangle$ and $\langle\omega_{oo^{\prime}}^{s}\rangle=\langle\rho_{o}\partial_{t}\rho_{o^{\prime}}+\rho_{o^{\prime}}\partial_{t}\rho_{o}\rangle$ respectively. The mean inferred EPR using the observable vorticities reads [88]:
@@
The inferred mean EP $\langle\dot{\Sigma}_{O^{\omega}}\rangle$ gives a lower bound on $\langle\dot{\Sigma}\rangle$. Integrating eq. 60 and using Jensen’s inequality leads to:
@@
Where, $\mathcal{C}_{oo^{\prime}}^{a}(t,\Delta t)=\rho_{o}(t+\Delta t)\rho_{o^{\prime}}(t)-\rho_{o^{\prime}}(t+\Delta t)\rho_{o}(t)$ and $\mathcal{C}_{oo^{\prime}}^{s}(t,\Delta t)=\rho_{o}(t+\Delta t)\rho_{o^{\prime}}(t)+\rho_{o^{\prime}}(t+\Delta t)\rho_{o}(t)$ characterizes the anti-symmetric and symmetric part of the observable state correlation. Here, $\Delta_{0}^{\tau}C_{oo^{\prime}}^{a}=C_{oo^{\prime}}^{a}(t_{i},\tau)-C_{oo^{\prime}}^{a}(t_{i},0)$ and $\Delta_{0}^{\tau}C_{oo^{\prime}}^{s}=C_{oo^{\prime}}^{s}(t_{i},\tau)-C_{oo^{\prime}}^{s}(t_{i},0)$ is the difference between the initial and final states.
Using $\{\omega_{oo^{\prime}}\}=\{\omega_{ij}\}$, the tightest possible bound on $\langle\dot{\Sigma}\rangle$ is obtained using state-space correlations. Physically, it signifies that considering the temporal cross-correlations between all linearly independent macrostates saturates the TUKR in the state space. The set of eqs. 60 and 61 is the formulation of short-time and finite-time TKUR using temporal correlations between the observable states. It should be compared to the quadratic [164] and non-quadratic [161, 165] counterparts.
@@ 5.1 Diffusive systems
We consider systems without reactive transitions; $\{\gamma\gamma^{\prime}\}=\emptyset$ is an empty set.
A Generalised Kawasaki-Dean equation
Considering $\vec{F}_{sp}=0$. The eq. 28 is reduced to:
@@ A Generalised Kawasaki-Dean equation
The eq. 62 generalizes the Kawasaki-Dean equation for the interacting systems [66, 67], valid in the small interaction parameter regime. In contrast eq. 62 encapsulates the effect of the interaction on the noise amplitude. For the ideal non-interacting reciprocal particle $\mu_{i}=\ln{\rho_{i}}$, one recovers the Kawasaki-Dean equation for the ideal particles.
@@ B Macroscopic fluctuation theory
Consider the systems without chemical reaction change between the particle types, the generalized macroscopic fluctuation theory EOM for it reads:
$\begin{split}\partial_{t}\rho_{i}&=\nabla^{\vec{\mathcal{D}}}\cdot\left\{D_{i}^{\vec{\mathcal{D}}}\left(\nabla^{\vec{\mathcal{D}}}\mu_{i}+\nabla^{\vec{\mathcal{D}}}F_{i}^{nr}-\vec{F}_{i}^{sp}\right)\right\}+\frac{1}{\sqrt{\Omega}}\nabla^{\vec{\mathcal{D}}}\cdot\left(\sqrt{{2D_{i}^{\vec{\mathcal{D}}}}}\>\vec{\hat{\xi}}_{i}\right).\end{split}$ (63)
The EOM eq. 63 the MFT formulation for non-reciprocal systems [87]. In contrast to Ref.[87], we have systematically derived the coarse-grained EOM with density-dependent mobility $D_{i}^{\vec{\mathcal{D}}}=d_{i}^{\vec{\mathcal{D}}}e^{\mu_{i}+F_{i}^{nr}}$. The second law of thermodynamics is $\langle\dot{\Sigma}\rangle=-d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])+\langle\dot{\Sigma}^{nr}\rangle+\langle\dot{\Sigma}^{sp}\rangle$. Here, $\langle\dot{\Sigma}^{sp}\rangle$ is a non-quadratic dissipation function in contrast to the quadratic one obtained in MFT [87].
@@ A Chemo sensing bacteria
We consider a prototypical system of bacteria(b) attracted toward a chemical(c). Thus, the microscopic interaction rules read $v_{bc}=-1$, $v_{cb}=0$, $v_{cc}=0$, $v_{bb}=0$. This leads to $v_{bc}^{r}=-\frac{1}{2}$ and $v_{cb}^{nr}=-v_{bc}^{nr}=\frac{1}{2}$. We choose this model with a minimal motif of non-reciprocity and its implications for diffusive dynamics. We consider the macroscopic fluctuating coarse-grained dynamics. The $E$ for it reads:
$\begin{split}E&=\int_{\mathcal{V}}\left[V_{bc}\rho c+\rho_{b}\ln{\left(\frac{\rho_{b}}{e}\right)}+\rho_{c}\ln{\left(\frac{\rho_{c}}{e}\right)}\right],\end{split}$ (64)
where, $V_{bc}^{r}=\beta\left(-\frac{1}{2}+\frac{\beta}{4\Omega}\right)$ and $V_{cb}^{nr}=\beta\left(\frac{1}{2}-\frac{\beta}{4\Omega}\right)$. Using $\mu_{b}^{r}=\log{(\rho)}+\rho_{c}V_{bc}^{r}$ and $\mu_{c}^{r}=\ln{(\rho_{c})}+\rho_{b}V_{bc}^{r}$ and $F_{b}^{nr}=-V_{cb}^{nr}\rho_{c}$ and $F_{c}^{nr}=\frac{1}{2}V_{cb}^{nr}\rho_{c}$. The mobility reads $D_{\rho}=\tilde{d}_{\rho}\rho e^{\rho_{c}(V_{bc}^{r}+V_{bc}^{nr})}=\tilde{d}_{\rho}\rho e^{-\beta\rho_{c}(1-\frac{\beta}{2\Omega})}$ and $D_{c}=\tilde{d}_{c}\rho_{c}$. $\vec{F}_{c}^{sp}=\vec{F}_{b}^{sp}=F_{bc}^{ch}=0$, thus, $\langle\dot{\Sigma}^{sp}\rangle=\langle\dot{\Sigma}^{ch}\rangle=0$. The expression for $\langle\dot{\Sigma}^{nr}\rangle$ reads:
@@
The second law of thermodynamics states $\langle\dot{\Sigma}\rangle=-d_{t}\langle E\rangle+d_{t}S^{gb}+\langle\dot{\Sigma}^{nr}\rangle$. Or equivalently, $\langle\dot{\Sigma}\rangle=-d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])+\langle\dot{\Sigma}^{nr}\rangle$. The $d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])$ quantifies the EPR due to the relaxation of $E$, which quantifies the relaxation of the symmetric macrostate correlations. The control parameters of the system are fixed thus $\langle W\rangle=0$.

### v3

@@ 4.4 Thermodynamic-Kinetic Uncertainty Relation
The thermodynamic uncertainty relation (TUR) delineates the trade-off between the precision of an observable current and the minimum EP required to sustain it [165, 240, 241, 242]. Here, the precision is the ratio between the square of the mean observable transition current and the variance of the transition current. The TUR gives a tighter bound on the EP than the second law of thermodynamics. In other words, the precision of an observable current comes at a minimum EP required to sustain it. The bounds on thermodynamic dissipation can be improved by choosing other appropriate physical observables and constraints [243, 244, 245, 246, 247, 248, 249, 250, 251, 252, 202, 244, 253, 254] and by its unification into the thermodynamic-kinetic uncertainty relation (TKUR) [255]. The TKUR’s connection to Speed Limits (SL) [256, 257, 258, 259, 260, 261, 262, 255, 172] and to the FR symmetry [263, 264, 265, 266, 267] has been established. Here, we demonstrate the TKUR for non-reciprocal systems, while a more fundamental analysis of the TKUR and non-quadratic speed limits is outlined in Refs. [89, 90].
@@ A Coarse-grained observable transition currents
We consider a set of macroscopic observable transition currents denoted by $O^{J}=\{\Delta^{o}\}$ and the corresponding mean inferred EPR $\langle\dot{\Sigma}_{O^{J}}\rangle$. The mapping $\{\{\Delta_{\gamma\gamma^{\prime}}\},\{\Delta_{i}^{\vec{\mathcal{D}}}\}\}=\{\Delta\}\to\{\Delta^{o}\}$ is considered to be many-to-one, so that each microscopic transition contributes to one macroscopic observable transition current. This mapping is represented using an observable matrix $\mathbb{O}^{J}$, with the row and column indices corresponding to the observable and microscopic currents, respectively. Mathematically, the mapping $\{\{\Delta_{\gamma\gamma^{\prime}}\},\{\Delta_{i}^{\vec{\mathcal{D}}}\}\}\to\{\Delta^{o}\}$ implies $\mathbb{O}_{o^{\prime}i^{\prime}}^{J}=1\implies\mathbb{O}_{o^{\prime\prime}i^{\prime}}^{J}=0,\forall o^{\prime\prime}\in\{\Delta^{o}\}-o^{\prime}$. The observable current and traffic vectors satisfy $\vec{J}_{\Delta^{o}}=\mathbb{O}^{J}\vec{J}_{\Delta}$ and $\vec{T}_{\Delta^{o}}=\mathbb{O}^{J}\vec{T}_{\Delta}$. In particular, using eqs. 30 and 32, one can equivalently obtain the same expressions for the mean and variance of the observable current. Then, the mean inferred EPR $\langle\dot{\Sigma}_{O^{J}}\rangle$ reads [89, 90]:
$\begin{split}\langle\dot{\Sigma}_{O^{J}}\rangle=\sum_{\{\Delta^{o}\}}2J_{\Delta^{o}}\tanh^{-1}{\left(\frac{J_{\Delta^{o}}}{T_{\Delta^{o}}}\right)}.\end{split}$ (59)
The inequality $\langle\dot{\Sigma}\rangle\geq\langle\dot{\Sigma}_{O^{J}}\rangle$ is proven using the generalized log-sum inequality and the definitions $\vec{J}_{\Delta^{o}}=\mathbb{O}^{J}\vec{J}_{\Delta}$ and $\vec{T}_{\Delta^{o}}=\mathbb{O}^{J}\vec{T}_{\Delta}$ [89, 90, 268]. The non-quadratic formulation gives a tighter bound on the EPR than the traditional quadratic TUR [262] and is closely related to [269, 260, 255, 172]. The mean observable inferred EP $\langle{\Sigma}_{O^{J}}\rangle$ satisfies the inequality [89, 90]:
@@
where $J_{o}^{\tau}=\int_{0}^{\tau}J_{o}dt$ and $T_{o}^{\tau}=\int_{0}^{\tau}T_{o}dt$ are the time-integrated current and traffic. We have utilized Jensen’s inequality to obtain the inequality in eq. 60 from the equality in eq. 59. The non-quadratic dissipation function gives a tighter bound than the quadratic TKUR, and its detailed physical and technical analysis is summarized in Refs.[89, 90].
@@ B Coarse-grained vorticities and state-correlations
Here, we aim to examine the EPR inference using state-space observables. We choose a specific sub-case of the set of vorticity currents between the observable macrostate $\rho_{o}=\mathbb{O}^{\rho}\vec{\rho}$, where $\mathbb{O}^{\rho}$ quantifies the many-to-one mapping from $\{\rho_{i}\}\to\{\rho_{o}\}$. $O^{\omega}=\{\omega_{oo^{\prime}}\}$ is quantified by the observable vorticity operator $\mathbb{O}^{\omega}$, which connects microscopic vorticities $\{\omega_{ij}\}$ to observable vorticities through $\vec{\omega}_{O}=\mathbb{O}^{\omega}\vec{\omega}$. The anti-symmetric (vorticity) and symmetric (traffic) components are denoted by $\langle\omega_{oo^{\prime}}\rangle=\langle\rho_{o}\partial_{t}\rho_{o^{\prime}}-\rho_{o^{\prime}}\partial_{t}\rho_{o}\rangle$ and $\langle\omega_{oo^{\prime}}^{s}\rangle=\langle\rho_{o}\partial_{t}\rho_{o^{\prime}}+\rho_{o^{\prime}}\partial_{t}\rho_{o}\rangle$, respectively. The mean inferred EPR using the observable vorticities reads [89, 90]:
@@
The inferred mean EP $\langle\dot{\Sigma}_{O^{\omega}}\rangle$ provides a lower bound on $\langle\dot{\Sigma}\rangle$. Integrating eq. 61 and using Jensen’s inequality leads to a bound on the inferred EP using the state correlations:
@@
Here, $\mathcal{C}_{oo^{\prime}}^{a}(t,\Delta t)=\rho_{o}(t+\Delta t)\rho_{o^{\prime}}(t)-\rho_{o^{\prime}}(t+\Delta t)\rho_{o}(t)$ and $\mathcal{C}_{oo^{\prime}}^{s}(t,\Delta t)=\rho_{o}(t+\Delta t)\rho_{o^{\prime}}(t)+\rho_{o^{\prime}}(t+\Delta t)\rho_{o}(t)$ characterize the anti-symmetric and symmetric parts of the observable state correlation. Furthermore, $\Delta_{0}^{\tau}C_{oo^{\prime}}^{a}=C_{oo^{\prime}}^{a}(t_{i},\tau)-C_{oo^{\prime}}^{a}(t_{i},0)$ and $\Delta_{0}^{\tau}C_{oo^{\prime}}^{s}=C_{oo^{\prime}}^{s}(t_{i},\tau)-C_{oo^{\prime}}^{s}(t_{i},0)$ denote the differences between the initial and final times.
Using $\{\omega_{oo^{\prime}}\}=\{\omega_{ij}\}$, the tightest possible bound on $\langle\dot{\Sigma}\rangle$ is obtained using state-space correlations. Physically, this signifies that considering the temporal cross-correlations between all linearly independent macrostates yields the tightest possible bound on the EP using the state-space TUKR. The set of eqs. 61 and 62 provides the formulation of the short-time and finite-time TKUR using temporal correlations between the observable states [89, 90]. This should be compared to the quadratic [170] and non-quadratic [167, 171] counterparts.
4.5 Information thermodynamics
The framework of information thermodynamics takes into account the thermodynamic implications of statistical information. Statistical information on a system’s state is encoded in Gibbs’s entropy eqs. 34 and 35. An information gain can be achieved through a feedback-controlled operation or measurement [270, 271, 236, 272, 273]. This information gain reduces the statistical uncertainty associated with the system, resulting in a lower EP, signifying the reduced statistical uncertainty.
We define the hidden variable as $Y$, which does not affect the transition dynamics (bulk term) of the relevant macrostates $\{\rho\}$. Then, the stochastic mutual information $I_{\{\rho\},Y}$ between them is defined as:
Here, Bayes’ theorem is used to reach the second equality in eq. 63. Mutual information quantifies the statistical correlation (or dependence) between $\{\rho\}$ and the hidden variable $Y$. It vanishes if the macrostate of the system is statistically independent of the hidden variable, that is, $\mathcal{P}_{\{\rho\}|Y}=\mathcal{P}_{\{\rho\}},\forall Y\in\mathcal{V}_{Y},\{\rho\}\in\mathcal{V}_{\{\rho\}}$. Here, $\mathcal{V}_{Y}$ and $\mathcal{V}_{\{\rho\}}$ denote the probability measure spaces for the hidden variable $Y$ and macrostates $\{\rho\}$, respectively. The mean mutual information is defined as $\langle I_{\{\rho\},Y}\rangle=\int_{\mathcal{V}_{Y}}\int_{\mathcal{V}_{\{\rho\}}}I_{\{\rho\},Y}\>\mathcal{P}_{\{\rho\},Y}$. Here, $\int\mathbb{D}[\{\rho\}]\equiv\int_{\mathcal{V}_{\{\rho\}}}$ is the shorthand notation used for summation over the probability measure space of the macrostate $\{\rho\}$. Then, the mutual information EPR is defined as $\langle\dot{I}_{\{\rho\},Y}\rangle=\int_{\mathcal{V}_{Y}}\int_{\mathcal{V}_{\{\rho\}}}I_{\{\rho\},Y}\>d_{t}\mathcal{P}_{\{\rho\},Y}$, whose exact expressions for the mutual information EP and EPR are:
$\begin{split}\langle I_{\{\rho\},Y}\rangle&=\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}|Y})-\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}}),\\ \langle\dot{I}_{\{\rho\},Y}\rangle&=d_{t}\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}|Y})-d_{t}\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}})\end{split}.$ (64)
Incorporating the mutual information EP and EPR, the second law of thermodynamics eq. 41 is modified. Thus, the Gibbs EPR is replaced by the conditional Gibbs EPR. The conditional Gibbs entropy satisfies the inequality $\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}})\geq\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}|Y})$. Physically, this corresponds to the information gained due to the hidden variable $Y$, which reduces the statistical ignorance (Gibbs entropy) of the macrostate $\{\rho\}$ by encapsulating its correlations with the hidden variable $Y$. $\langle\dot{I}_{\{\rho\},Y}\rangle\leq 0$ holds, which gives the sign convention for the information gain. Equation 64 satisfies the equality, $\langle\dot{I}_{\{\rho\},Y}\rangle=d_{t}\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\},Y})-d_{t}\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}})-d_{t}\mathcal{S}^{gb}(\mathcal{P}_{Y})$. Similarly, the fluctuation theorem follows trivially by modifying eq. 55 to incorporate stochastic mutual information eq. 63, or equivalently, by replacing the reference state entropy with the reference conditional state entropy $S_{E}^{state}\to S_{E}^{state|Y}=-\log{(\mathcal{P}_{\{\rho\}|Y}/\mathcal{P}^{E}_{\{\rho\}|Y})}$. This is a manifestation of the physical symmetry of measurement (certainty) and uncertainty. We have briefly highlighted the procedure to incorporate information thermodynamics for non-reciprocal systems. The rest is rather trivial.
@@ 5.1 Diffusive systems
We consider systems without reactive transitions, for which $\{\gamma\gamma^{\prime}\}=\emptyset$ is an empty set.
A Generalized Kawasaki-Dean equation
Considering $\vec{F}_{sp}=0$, eq. 29 is reduced to [87]:
@@ A Generalized Kawasaki-Dean equation
Equation 65 generalizes the Kawasaki-Dean equation for interacting systems [66, 67], valid in the regime of small interaction parameters. In contrast, eq. 65 encapsulates the effect of the interaction on the amplitude of the noise. For an ideal non-interacting reciprocal particle, $\mu_{i}=\ln{\rho_{i}}$, one recovers the Kawasaki-Dean equation for the ideal particles.
@@ B Macroscopic fluctuation theory
We consider systems without chemical reaction changes between the particle types. In this case, the generalized macroscopic fluctuation theory EOM for the macrostate reads:
$\begin{split}\partial_{t}\rho_{i}&=\nabla^{\vec{\mathcal{D}}}\cdot\left\{D_{i}^{\vec{\mathcal{D}}}\left(\nabla^{\vec{\mathcal{D}}}\mu_{i}+\nabla^{\vec{\mathcal{D}}}F_{i}^{nr}-\vec{F}_{i}^{sp}\right)\right\}\\ &\hskip 28.45274pt+\frac{1}{\sqrt{\Omega}}\nabla^{\vec{\mathcal{D}}}\cdot\left(\sqrt{{2D_{i}^{\vec{\mathcal{D}}}}}\>\vec{\hat{\xi}}_{i}\right).\end{split}$ (66)
Equation 66 is the MFT formulation for non-reciprocal systems [88]. In contrast to Ref.[88], we have systematically derived the coarse-grained EOM with the density-dependent mobility $D_{i}^{\vec{\mathcal{D}}}=d_{i}^{\vec{\mathcal{D}}}e^{\mu_{i}+F_{i}^{nr}}$ [87]. The second law of thermodynamics for this system is $\langle\dot{\Sigma}\rangle=-d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])+\langle\dot{\Sigma}^{nr}\rangle+\langle\dot{\Sigma}^{sp}\rangle$. Here, $\langle\dot{\Sigma}^{sp}\rangle$ is a non-quadratic dissipation function, in contrast to the quadratic one obtained in MFT [88], therefore avoiding the underestimation of the dissipation for far-from-equilibrium systems [89, 90].
@@ A Chemo sensing bacteria
We consider a prototypical system of bacteria (b) attracted toward a chemical (c). Thus, the microscopic interaction rules for this physical system read $v_{bc}=-1$, $v_{cb}=0$, $v_{cc}=0$, $v_{bb}=0$. This leads to $v_{bc}^{r}=-\frac{1}{2}$ and $v_{cb}^{nr}=-v_{bc}^{nr}=\frac{1}{2}$. We choose this model with a minimal motif of non-reciprocity and its implications for diffusive dynamics; therefore, the self-propulsion forces vanish, $\vec{f}_{c}^{sp}=\vec{f}_{b}^{sp}=0$. The macroscopic energy functional $E$ for the density of the bacteria ($\rho_{b}$) and the chemical ($\rho_{c}$) reads:
$\begin{split}E&=\int_{\mathcal{V}}\left[V_{bc}\rho_{b}\rho_{c}+\rho_{b}\ln{\left(\frac{\rho_{b}}{e}\right)}+\rho_{c}\ln{\left(\frac{\rho_{c}}{e}\right)}\right],\end{split}$ (67)
where $V_{bc}^{r}=\beta\left(-\frac{1}{2}+\frac{\beta}{4\Omega}\right)$ and $V_{cb}^{nr}=\beta\left(\frac{1}{2}-\frac{\beta}{4\Omega}\right)$. Using $\mu_{b}^{r}=\log{(\rho_{b})}+\rho_{c}V_{bc}^{r}$ and $\mu_{c}^{r}=\ln{(\rho_{c})}+\rho_{b}V_{bc}^{r}$, and $F_{b}^{nr}=-V_{cb}^{nr}\rho_{c}$ and $F_{c}^{nr}=\frac{1}{2}V_{cb}^{nr}\rho_{c}$. The mobility of the bacterial density and chemical density are $D_{\rho}=\tilde{d}_{\rho}\rho_{b}e^{\rho_{c}(V_{bc}^{r}+V_{bc}^{nr})}=\tilde{d}_{\rho}\rho_{b}e^{-\beta\rho_{c}(1-\frac{\beta}{2\Omega})}$ and $D_{c}=\tilde{d}_{c}\rho_{c}$, respectively. Thus, $\langle\dot{\Sigma}^{sp}\rangle=\langle\dot{\Sigma}^{ch}\rangle=0$. The expression for $\langle\dot{\Sigma}^{nr}\rangle$, which quantifies the thermodynamic cost of maintaining the vorticity currents between the bacteria and the chemical, reads:
@@
The second law of thermodynamics states $\langle\dot{\Sigma}\rangle=-d_{t}\langle E\rangle+d_{t}S^{gb}+\langle\dot{\Sigma}^{nr}\rangle$. Or equivalently, $\langle\dot{\Sigma}\rangle=-d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])+\langle\dot{\Sigma}^{nr}\rangle$. Here, $d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])$ quantifies the EPR due to the relaxation of $E$, which equivalently quantifies the relaxation of the symmetric macrostate correlations. $\langle W\rangle=0$, since the control parameters of the system are fixed.